OpenStudy (anonymous):

I will give you a medal and be your fan Greta had 5/8 pounds of ham.She used 1/3 pound for sandwhiches. How many pounds of ham does she have left? A. 5/24ib B. 7/24 ib C. 1/3 ib D. 4/5 ib

OpenStudy (calculusxy):

subtract 5/8 - 1/3 to get your answer :)

OpenStudy (anonymous):

4/5? You're not suppose to subtract the demonimators are you?

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

\[\frac{5}{8}-\frac{1}{3}=\frac{5\times 3-8\times 1}{8\times 3}\]

OpenStudy (anonymous):

that is how you subtract fractions \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\]

OpenStudy (a1234):

5/8 - 1/3 Make them a COMMON denominator: 15/24 - 8/24 15 - 8 = 7 7/24

OpenStudy (anonymous):

Thank you guys for the help I understand :)